### v1

# Prompt-based mental health screening from social media text

Abstract
This article presents a method for prompt-based mental health screening from a large and noisy dataset of social media text. Our method uses GPT 3.5. prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. Results are found to be on pair with a BERT mixture of experts classifier, and incurring only a fraction of its computational costs.
1 Introduction
Textual representation models used in NLP applications have changed dramatically in recent years, moving on from simple token counts (e.g., bag-of-words) to transformer-based and large language models (LLMs) such as GPT [^1] https://platform.openai.com/docs/models and Bloom [^2] https://huggingface.co/bigscience/bloom, among many others. In particular, LLMs have successfully circumvented the need for labelled training data entirely, in so-called prompt-based methods that are now mainstream in the field, and have been shown to obtain higher accuracy in a wide range of tasks.
Despite its potential to reduce training costs, however, LLM prompting still requires careful consideration in applications involving large or noisy data, as it is often the case of social media. In particular, we notice that social media data may be available as a long history of publications, often stretching over thousands of posts, and yet many or most may be unrelated to the underlying task. This may be the case, for instance, when screening for user-level information such as mental health statuses (e.g., related to depression), a task that may challenge supervised and prompt-based methods alike.
Based on these observations, this work investigates the use of LLM prompting as an aid to mental health screening from social media text. Taking the issue of depression detection from Brazilian Twitter timelines as a case study, we propose a method that combines GPT prompting and standard bag-of-words classification for fast, computationally inexpensive results, and which is evaluated against SOTA results obtained by BERT mixture of experts [dos Santos et al. 2023b].
2 Related work
Depression detection may be seen as an instance of author profiling [Rangel et al. 2020, Flores et al. 2022, Pavan et al. 2023] from social media text. Recent studies in the field are summarised in Table 1, categorised by text genre (Reddit, Twitter), text model (b=bag of words, BERT, e=embeddings, s=sentiment), and methods.
Table 1: Depression prediction from text data.
Study Genre Text model Methods
[Cohan et al. 2018] reddit b,e FastText
[Aragón et al. 2019] reddit s SVM
[Burdisso et al. 2020] reddit b SS3
[Lin et al. 2020] twitter e CNN
[Souza et al. 2020] reddit e LSTM
[Souza et al. 2021] reddit e LSTM+CNN
[Ansari and Ji 2022] reddit,twitter e,s LR+LSTM
[dos Santos et al. 2023a] twitter BERT Bi-LSTM
[dos Santos et al. 2023b] twitter BERT mixture of experts
We notice a slight predominance of works based on Reddit. This may be explained by the greater ease of access and reuse of this type of data for research purposes, which is far more restricted in the case of Twitter/X. Reddit publications are taken as the basis for some of the best-known datasets available for the English language, including the SMHD [Cohan et al. 2018] and eRisk [Parapar et al. 2022] corpora.
As for the kinds of textual model under consideration, Table 1 reflects the natural evolution of the NLP and related fields, with an initial prevalence of bag-of-words models and feature engineering, and its gradual replacement by models based on word embeddings and, more recently, transformers. With regard to the computational methods used, a similar trajectory is generally observable, with the use of traditional (e.g., linear) classifiers being gradually replaced by sequence classification methods based on deep learning, including the more recent use of transformer-based architectures. In both cases - text representation and methods - we notice also that better results are usually accompanied by an increase in computational costs.
Finally, we notice that, with the exception of the work based on the SetembroBR [dos Santos et al. 2023a, dos Santos et al. 2020] corpus to be used in the present work, all of the above studies are devoted to the English language, and none make use of prompt-based methods to query a LLM for depression directly.
3 Method
As in [Cohan et al. 2018, Parapar et al. 2022] and others, our approach to depression detection relies on a dataset of social media (in our case, Twitter/X) texts that have been published by either individuals who self-reported a depression diagnosis, or by a random (control) group. Thus, the task at hand represents a binary classification problem intended to distinguish individuals who will most likely receive a depression diagnosis (called Diagnosed class) in the future from the general population (called Control class), and not to distinguish depressed from non-depressed individuals per se.
3.1 Data
We use of the depression portion of the SetembroBR corpus [dos Santos et al. 2023a], a collection of Twitter timelines published by Diagnosed and Control individuals in which only the data prior to the diagnosis date is kept. As in other language resources of this kind, the Control (i.e., random) subset is designed so as to be seven times larger than the Diagnosed group, making a heavily imbalanced classification task. Table 2 presents data descriptive statistics.
Table 2: SetembroBR depression corpus descriptive statistics.
Statistics Diagnosed Control Overall
Users (timelines) 1,684 11,788 13,472
Words (million) 29.32 201.94 231,26
Publications (million) 2.43 16.99 19,42
3.2 Approach
Screening for depression from social media gives rise to the questions of how to handle a large number of publications, most of which unlikely to be relevant to the task. To this end, we envisaged a prompt-based approach called Prompt.Bow that relies on GPT 3.5 to enrich an otherwise standard text classifier.
First, GPT 3.5. prompting is used to assess every tweet in our data according to relevance for mental health prediction. The (English-translated) prompt in question, adapted from the clinical description of depression in [American Psychiatric Association 2013], is shown in Figure 1.
Figure 1: Prompt instruction to GPT.
By submitting this prompt to each of the 19.42 million tweets in the corpus, the data was categorised as having high (1), medium (2) or low (3) relevance to mental health. Table shows the label distribution across the training portion of the corpus.
Table 3: Training data distribution according to relevance for depression.
Diagnosed class Control class
Relevance Tweets % Tokens % Tweets % Tokens %
high 39k 2.0% 674k 5.4% 153k 1.1% 2,372k 3.0%
medium 180k 9.3% 3,663k 29.4% 1,031k 7.6% 20,566k 25.7%
low 1,727k 88.7% 8,122k 65.2% 12,439k 91.3% 56,971k 71.3%
As expected, most publications (65.2% in the Diagnosed class, and 71.3% in the Control class) are deemed of low relevance for mental health prediction according to the prompted model. On the other hand, highly relevant publications are relatively rare (5.4% in the Diagnosed class, and 3.0% in the Control class).
More importantly, however, when prompting an LLM for mental health we are largely focusing on semantics, that is, on symptoms and other well-known clinical signs of depression, but this is not to say that other factors can or should be overlooked. For instance, our current prompt allows the model to pinpoint a wide range of publications that may suggest, e.g., eating disorders or negative language use, both of which known to be related to depression, cf. [American Psychiatric Association 2013], but it does not explicitly account more fine-grained linguistic indicators of depression such as the use of first person pronouns [Trifu et al. 2017], absolute terms [Al-Mosaiwi and Johnstone 2018] and others. These indicators, which may be present in any publication of low or high relevance to depression alike, are also important predictors of mental health statuses, and no data point should in principle be discarded solely based on the LLM output.
As a means to keep the full train data available to the classifier while distinguishing between more and less relevant messages (which clearly show different distributions across classes in Table 3), Prompt.Bow uses the category labels provided by the LLM to split the training data into high, medium, and low relevance subsets, from which we create three individual bag-of-words vectors to be concatenated as discussed below.
Each of the three vectors consists of TF-IDF counts of texts labelled as high/medium/low by the GPT prompt method, and are further reduced using univariate feature selection using F1 as a score function. The final ‘high’ and ‘medium’ vectors were reduced to $k=$ 6000 features each, and the final ‘low’ vector was reduced to $k=$ 3000 features, all of which concatenated as a single 15000-word vector.
In addition to distinguishing between high-, medium- and low-relevance messages, we use the information provided by the LLM ALSO to HELP capture message order, the underlying assumption being that certain patterns (e.g., a series of consecutive ‘highly relevant’ messages) may be indicative of depression. To this end, we created a bigram model of high/medium/low labels only, reduced to its $k=$ 40000 most relevant bigram features by performing F1 univariate feature selection, and then appended to the previous 15000-word vector.
The resulting vector is taken as the input to a standard logistic regression classifier. This choice is motivated by the observation that much of the deep (e.g., semantics-driven) language processing had already been performed by the LLM, and that a simple text classifier should suffice for user label prediction.
4 Evaluation
Table 4 presents the results of the present Prompt.Bow model alongside the results reported in [dos Santos et al. 2023b] for BERT mixture of experts using BERTabaporu [da Costa et al. 2023], which currently stands as the SOTA for the present setting. In both cases, results are based on the standard test portion of the corpus.
Table 4: Classification results.
Model Precision Recall F1
Prompt.BoW 0.64 0.72 0.66
BERT.MoE 0.64 0.67 0.65
Table 4 shows that results remain close, with a small advantage for the Prompt.Bow approach over BERT mixture of experts. However, by relying on an input provided by the pre-trained LLM, we notice that these results were obtained by using a computationally inexpensive classifier model, which represents a stark contrasts to computationally-intensive BERT mixture of experts supervision.
5 Final remarks
This work presented an experiment in prompt-based mental health screening from a large, noisy corpus of social media publications that relies on LLM prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. This was shown to obtain competitive results if compared to a BERT-based classifier architecture that represents the SOTA in the present setting, but incurring only a fraction of its computational costs.
As future work, we intend to refine the present method by using the formal definition of depression in [American Psychiatric Association 2013] as a prompt, and by fully integrating symptoms and linguistic indicators within a neural architecture for both depression and anxiety disorder prediction.
6 Acknowledgements
The present work has been financed by the São Paulo Research Foundation (FAPESP grant #2021/08213-0). The first author has been supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 - grant # 88887.475847/2020-00. This work was carried out at the Center for Artificial Intelligence (C4AI-USP), with support by the São Paulo Research Foundation (FAPESP grant #2019/07665-4) and by the IBM Corporation.

### v2

@@
 prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels.
-Results are found to be on pair with a BERT mixture of experts classifier, and incurring only a fraction of its computational costs.
+Results are found to be on pair with a BERT mixture of experts classifier, and incurring only a fraction of its training costs.
 1 Introduction
-Textual representation models used in NLP applications have changed dramatically in recent years, moving on from simple token counts (e.g., bag-of-words) to transformer-based and large language models (LLMs) such as GPT [^1] https://platform.openai.com/docs/models and Bloom [^2] https://huggingface.co/bigscience/bloom, among many others.
+Textual representation models used in NLP applications have changed dramatically in recent years, moving on from simple token counts (e.g., bag-of-words) [Pereira and Paraboni 2007] to transformer-based and large language models (LLMs) such as GPT [^1] https://platform.openai.com/docs/models and Bloom [^2] https://huggingface.co/bigscience/bloom, among many others.
 In particular, LLMs have successfully circumvented the need for labelled training data entirely, in so-called prompt-based methods that are now mainstream in the field, and have been shown to obtain higher accuracy in a wide range of tasks.
@@
 2022, Pavan et al.
-2023] from social media text.
+2023, Pavan et al.
+2020] from social media text.
 Recent studies in the field are summarised in Table 1, categorised by text genre (Reddit, Twitter), text model (b=bag of words, BERT, e=embeddings, s=sentiment), and methods.
@@
 2023a, dos Santos et al.
-2020] corpus to be used in the present work, all of the above studies are devoted to the English language, and none make use of prompt-based methods to query a LLM for depression directly.
+2020] corpus to be used in the present work, all of the above studies are devoted to the English language and, to the best of our knowledge, none make use of prompt-based methods to query a LLM for depression directly.
 3 Method
@@
 2022] and others, our approach to depression detection relies on a dataset of social media (in our case, Twitter/X) texts that have been published by either individuals who self-reported a depression diagnosis, or by a random (control) group.
-Thus, the task at hand represents a binary classification problem intended to distinguish individuals who will most likely receive a depression diagnosis (called Diagnosed class) in the future from the general population (called Control class), and not to distinguish depressed from non-depressed individuals per se.
+Thus, the task at hand represents a binary classification problem intended to distinguish individuals who will most likely receive a depression diagnosis (called Diagnosed class) in the future from the general population (called Control class), and not to distinguish depressed from non-depressed individuals per se [^3] In fact, the data does not convey any guaranteed ‘non-depressed’ individuals [dos Santos et al.
+2023a]..
 3.1 Data
@@
 Publications (million) 2.43 16.99 19,42
+In the present work, we follow the standard train/test split provided by the corpus, as described in [dos Santos et al. 2023a].
 3.2 Approach
@@
 To this end, we envisaged a prompt-based approach called Prompt.Bow that relies on GPT 3.5 to enrich an otherwise standard text classifier.
-First, GPT 3.5.
-prompting is used to assess every tweet in our data according to relevance for mental health prediction.
+These two steps - prompting and classification - are described individually as follows.
+First as a means to identify messages that are potentially related to mental health, we use GPT 3.5.
+prompting to assess a random sample of 30,000 tweets.
 The (English-translated) prompt in question, adapted from the clinical description of depression in [American Psychiatric Association 2013], is shown in Figure 1.
 Figure 1: Prompt instruction to GPT.
-By submitting this prompt to each of the 19.42 million tweets in the corpus, the data was categorised as having high (1), medium (2) or low (3) relevance to mental health.
-Table shows the label distribution across the training portion of the corpus.
-Table 3: Training data distribution according to relevance for depression.
+By submitting this prompt to each of the 30,000 sample tweets, the data was categorised as having high (1), medium (2) or low (3) relevance to mental health.
+These GPT-labelled data was then taken as an input to train a T5 classifier [Raffel et al.
+2020] to label the entire 19.42-million tweets in the corpus.
+Table shows the label distribution across the training portion of the data.
+Table 3: Training data distribution according to relevance for depression. Tweets and tokens are shown in thousand units.
 Diagnosed class Control class
 Relevance Tweets % Tokens % Tweets % Tokens %
-high 39k 2.0% 674k 5.4% 153k 1.1% 2,372k 3.0%
-medium 180k 9.3% 3,663k 29.4% 1,031k 7.6% 20,566k 25.7%
-low 1,727k 88.7% 8,122k 65.2% 12,439k 91.3% 56,971k 71.3%
+high 39 2.0% 674 5.4% 153 1.1% 2,372 3.0%
+medium 180 9.3% 3,663 29.4% 1,031 7.6% 20,566 25.7%
+low 1,727 88.7% 8,122 65.2% 12,439 91.3% 56,971 71.3%
 As expected, most publications (65.2% in the Diagnosed class, and 71.3% in the Control class) are deemed of low relevance for mental health prediction according to the prompted model.
 On the other hand, highly relevant publications are relatively rare (5.4% in the Diagnosed class, and 3.0% in the Control class).
-More importantly, however, when prompting an LLM for mental health we are largely focusing on semantics, that is, on symptoms and other well-known clinical signs of depression, but this is not to say that other factors can or should be overlooked.
+More importantly, however, when prompting an LLM for mental health we are largely focusing on semantics, that is, on symptoms and other well-known clinical signs of depression.
 For instance, our current prompt allows the model to pinpoint a wide range of publications that may suggest, e.g., eating disorders or negative language use, both of which known to be related to depression, cf.
-[American Psychiatric Association 2013], but it does not explicitly account more fine-grained linguistic indicators of depression such as the use of first person pronouns [Trifu et al.
-2017], absolute terms [Al-Mosaiwi and Johnstone 2018] and others.
+[American Psychiatric Association 2013], but this is not to say that other factors can or should be overlooked.
+In particular, we notice that LLM prompting may not explicitly account for more fine-grained linguistic indicators of depression such as the use of first person pronouns [Trifu et al.
+2017], absolute terms [Al-Mosaiwi and Johnstone 2018] and other lexical factors (e.g., denoting emotion as in hate speech, cf.
+[da Silva et al.
+2020]).
 These indicators, which may be present in any publication of low or high relevance to depression alike, are also important predictors of mental health statuses, and no data point should in principle be discarded solely based on the LLM output.
-As a means to keep the full train data available to the classifier while distinguishing between more and less relevant messages (which clearly show different distributions across classes in Table 3), Prompt.Bow uses the category labels provided by the LLM to split the training data into high, medium, and low relevance subsets, from which we create three individual bag-of-words vectors to be concatenated as discussed below.
-Each of the three vectors consists of TF-IDF counts of texts labelled as high/medium/low by the GPT prompt method, and are further reduced using univariate feature selection using F1 as a score function.
-The final ‘high’ and ‘medium’ vectors were reduced to $k=$ 6000 features each, and the final ‘low’ vector was reduced to $k=$ 3000 features, all of which concatenated as a single 15000-word vector.
-In addition to distinguishing between high-, medium- and low-relevance messages, we use the information provided by the LLM ALSO to HELP capture message order, the underlying assumption being that certain patterns (e.g., a series of consecutive ‘highly relevant’ messages) may be indicative of depression.
-To this end, we created a bigram model of high/medium/low labels only, reduced to its $k=$ 40000 most relevant bigram features by performing F1 univariate feature selection, and then appended to the previous 15000-word vector.
-The resulting vector is taken as the input to a standard logistic regression classifier.
+As a means to keep the full train data available to the classifier whilst distinguishing between more and less relevant messages (which clearly show different distributions across classes in Table 3), Prompt.Bow uses the category labels provided by the LLM to split the training data into high, medium, and low relevance subsets, from which we create three individual bag-of-words vectors (to be concatenated as discussed below).
+In other words, the training texts are split into three (low/medium/high relevance) categories according to the previous GPT prompt method, and we build an individual BoW model from each of these three subsets.
+The resulting vectors are further reduced using univariate feature selection using F1 as a score function.
+The final ‘high’ and ‘medium’ vectors were reduced to $k=$ 6,000 features each, and the final ‘low’ vector was reduced to $k=$ 3,000 features, all of which concatenated as a single 15,000-word vector.
+In addition to distinguishing between high-, medium- and low-relevance messages in this way, we use the information provided by the LLM also to help capture message order, the underlying assumption being that certain patterns (e.g., a series of consecutive ‘highly relevant’ messages) may be indicative of depression.
+To this end, we created a bigram model of high/medium/low labels only, that is, representing sequences of ‘high’, ‘medium’ and ‘low’ labels only (and not text).
+This was further reduced to its $k=$ 40,000 most relevant (bigram) features by performing F1 univariate feature selection, and then appended to the previous 15,000-word vector.
+The resulting vector - a combination of three text models with different degrees of relevance to mental health, and a model that captures sequences of relevant messages - is taken as the input to a standard logistic regression classifier.
 This choice is motivated by the observation that much of the deep (e.g., semantics-driven) language processing had already been performed by the LLM, and that a simple text classifier should suffice for user label prediction.
@@
 Table 4 shows that results remain close, with a small advantage for the Prompt.Bow approach over BERT mixture of experts.
-However, by relying on an input provided by the pre-trained LLM, we notice that these results were obtained by using a computationally inexpensive classifier model, which represents a stark contrasts to computationally-intensive BERT mixture of experts supervision.
+However, by relying on an input provided by the pre-trained LLM, we notice that these results were obtained by using a computationally inexpensive classifier model (that is, leaving aside the costs of pre-training the LLM in the first place), which represents a stark contrasts to computationally-intensive BERT mixture of experts supervision.
 5 Final remarks
 This work presented an experiment in prompt-based mental health screening from a large, noisy corpus of social media publications that relies on LLM prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels.
-This was shown to obtain competitive results if compared to a BERT-based classifier architecture that represents the SOTA in the present setting, but incurring only a fraction of its computational costs.
+This was shown to obtain competitive results if compared to a BERT-based classifier architecture that represents the SOTA in the present setting, but incurring only a fraction of its training costs.
 As future work, we intend to refine the present method by using the formal definition of depression in [American Psychiatric Association 2013] as a prompt, and by fully integrating symptoms and linguistic indicators within a neural architecture for both depression and anxiety disorder prediction.